1)    An 8-bit register contains the Hex value A9. What is the register value after an arithmetic shift right? Starting from the initial number A9, determine the register value

 i)        after logical shift right



1
Expert's answer
2021-09-27T01:44:15-0400

A hex number, short for hexadecimal number, is a number expressed in the hexadecimal positional numeral system with a base of 16 which uses sixteen symbols: the numbers from 0 to 9 and the letters A,B,C,D,E, and F, to represent values from 0 to 15. Equivalently, small-case letters a through f can be used. For example, 10 in decimal is A in hex, 100 in decimal is 64 in hex, while 1,000 in decimal is 3E8 in hex. Hex numbers have signs, just like decimal ones, for example -1e is equal to -30 in decimal.

Hex numerals are used mostly in computing: computer system designerssoftware engineers and programmers as a convenient representation of the underlying binary systems. People in such occupations are most likely to be in need of a hex calculator or hex converter.

An ordinary person would encounter them every day when browsing the web since in website addresses (URLs) special characters are encoded as a hex numeral, e.g. %20 stands for "space" (blank). A lot of webpages also encode special characters in html in terms of their hexadecimal numerical character reference (&#xcode), e.g. ’ is the Unicode for a single quotation mark (’). A regular person browsing the web should not require a hexadecimal converter or calculator to do so.

Each position in a hexadecimal numeral represents a power of 16 the same way each position in a decimal number represents a power of 10. For example, the number 20 in decimal is 2 · 101 + 0 · 100 = 20. The hex number 20 is then 2 · 161 + 0 · 160 = 32 in dec. Similarly, the number 1E is 1 · 16 + 14 · 1 = 30 in decimal.

The process of hex to decimal conversion is therefore to first take each position and convert it to decimal, e.g. 9 is 9, but B is converted to 11, then to multiply each position by 16 to the power of the position number, counting from right to left and starting at zero. If you need to calculate large exponents like 168 you might find our exponent calculator useful.